# Bell Nonlocality from Wigner Negativity in Qudit Stabilizer States

### v3

@@ Abstract
Nonlocality is an essential concept that distinguishes quantum from classical models and has been extensively studied in systems of qubits. For higher-dimensional systems, certain results for their two-level counterpart, like Bell violations with stabilizer states and Clifford operators, do not generalize. On the other hand, similar to continuous variable systems, Wigner negativity is necessary for nonlocality in qudit systems. We show how a bipartite entangled qudit state can serve as a witness for nonlocality when it exhibits Wigner negativity. Additionally, we propose a new generalization of the CHSH inequality for qudits by inquiring correlations related to the Wigner negativity of stabilizer states under the adjoint action of a generalization of the qubit $\pi/8$-gate. A specified stabilizer state maximally violates the inequality among all qudit states based on its Wigner negativity. The Bell operator not only serves as a measure for the singlet fraction but also quantifies the volume of Wigner negativity. Furthermore, we give deterministic Bell violations, as well as violations with a constant number of measurements, for the Bell state relying on operators innate to higher-dimensional systems than the qudit at hand.
@@
This work focuses on qudit stabilizer states, which are uniquely defined by the Abelian subgroup of the qudit Pauli operators or Heisenberg-Weyl displacement operators, and encompass Bell states, GHZ states, and graph states. Contrary to qubits, Pauli operators on stabilizer states cannot invoke contextual violations —a prerequisite to nonlocality— due to their semi-classical nature exposed by a positive semi-definite Wigner function Kochen1975 , Howard2013 , Gross2006 . Similar to continuous variable systems PhysRevLett.129.230401 , the negativity of Gross’s Wigner function is equivalent to the contextuality of projective Pauli measurements in odd dimensions Howard2014 , Delfosse2017 . Therefore, Wigner-negativity, whether located in the state or the measurement, is a necessary requirement for nonlocality.
@@
For any odd prime dimension $d$, we first show in Eq. (11) how certain bipartite entangled states that exhibit Wigner negativity can be witnesses for nonlocality. Then, we construct a family of Bell operators on two qudits in Eq. (16) that relates to Gross’s Wigner function and provides a measure for the singlet fraction as well as for the volume of Wigner negativity. The Bell operators consist of qudit Pauli operators rotated by a generalization of the qubit $T$- or $\pi/8$-gate introduced in Howard2012 , which we call unitary cube operators. Mapping the Pauli group to the qudit Clifford group appleby2005symmetric , Appleby2009 , unitary cube operators cause Wigner negativity in stabilizer states when these states are rotated by them. A corresponding Bell inequality is violated by the Bell state and is a natural extension of the Clauser-Horne-Shimony-Holt (CHSH) inequality for qubits PhysRevLett.23.880 , which is similarly a linear combination of Pauli operators under the local rotation of the $T$-gate. We can then reduce the number of measurement settings constituting the Bell operator by only testing the stabilizer elements of the Bell state with the same implications as before. While the underlying mechanisms for Bell violations are equivalent to the ones Howard2015bell , our main contribution is a broader generalization of the qubit CHSH inequality for qudits adaptable to operators beyond the unitary cube operators, as well as a general insight into the origins of nonlocality in qudit systems. Lastly, this Bell operator can be adapted to any entangled qudit stabilizer state that maximally violates a corresponding Bell inequality, though it does not detect genuine multipartite entanglement beyond three parties. For $d>3$, unitary cube operators are inherently defined in the qudit system, that is, their eigenvalues are $d^{\text{th}}$ roots of unity, called characters of the finite field wiki:Dirichlet_character . Alternatively, we show strong nonlocal violations for Bell states in $d$ dimension with operators whose spectrum contains characters of yet higher-dimensional systems.
Setting.
@@
The prime setting for nonlocality, in a Bell scenario for $n$ parties, each party performs a local measurement $O_{m_{i}}$ depending on the local measurement setting $m_{i}$. Physical models can pose restrictions on the general probability $p({\bm{a}}|{\bm{m}})=p((a_{i})_{i}|(m_{i})_{i})$ of the outcome ${\bm{a}}$ for the measurement settings ${\bm{m}}$. In the quantum model, the probability distribution is derived from Born’s rule $p({\bm{a}}|{\bm{m}})=\mathrm{Tr}\left(\Pi^{{\bm{m}}}_{a^{\ast}}\rho\right)$ for a $n$-partite quantum state $\rho$ and projectors $\Pi^{{\bm{m}}}_{a^{\ast}}$ of the joint measurement $\bigotimes_{i=1}^{n}O_{m_{i}}$ with outcome $a^{\ast}=\prod_{i=1}^{n}a_{i}$. The (classical) local-hidden-variable (lhv) model is more restrictive with probability distributions of the form $p({\bm{a}}|{\bm{m}})=\sum_{\lambda}\mu(\lambda)\prod_{i=1}^{n}p_{i}(a_{i}|m_{i},\lambda)\,,$ for probability distributions $\mu$ and $p_{i}$. The separation can be proven with a Bell inequality for a Bell operator $\mathcal{B}=\sum_{{\bm{m}}}c_{{\bm{m}}}\bigotimes_{i=1}^{n}O_{m_{i}}\,,$ with coefficients $c_{{\bm{m}}}$. If an lhv model describes the probability distribution $p$, then
@@
and $\left\langle\mathcal{B}\right\rangle_{\mathrm{lhv}}\leq\max_{p\in\mathrm{lhv}}\left\langle\mathcal{B}\right\rangle_{\mathrm{lhv}}:=B_{\mathrm{lhv}}$ is a Bell inequality. Any distribution $p({\bm{a}}|{\bm{m}})$ such as $\sum_{{\bm{a}},{\bm{m}}}c_{{\bm{m}}}{\bm{a}}^{\ast}p({\bm{a}}|{\bm{m}})>B_{\mathrm{lhv}}$ defies classical local models. In this work, the measurements that define the Bell experiment are projective measurements on qudits with non-degenerate eigenvalues as outcomes.
To prevent studying properties inherent to subsystems, which is contiguous in systems of composite dimensions, we restrict our study to qudit systems of prime dimension $d$. For example, consider the $n$-party GHZ state $|\text{GHZ}^{d}\rangle=\sum^{d-1}_{k=0}|k\rangle^{\otimes n}/\sqrt{d}$. If $d=pq$ for two integers $q,p$, the GHZ state is a product state over two composite systems with dimensions $p$ and $q$, $|\text{GHZ}^{p}\rangle\otimes|\text{GHZ}^{q}\rangle=\sum^{p-1}_{j=0}|j\rangle^{\otimes n}\otimes\sum^{q-1}_{k=0}|k\rangle^{\otimes n}/\sqrt{pq}=\sum^{p-1}_{j=0}\sum^{q-1}_{k=0}|j\otimes k\rangle^{\otimes n}/\sqrt{d}=|\text{GHZ}^{d}\rangle$ with the representation of $j\otimes k\sim k+jp$. As the case $d=2$ corresponds to qubits, this leaves us with odd prime dimension, where all numbers are integers modulo $d$ forming a finite field, in particular $2^{-1}$ is unambiguous.
Pauli group and Wigner function.
The qudit operators $X$ and $Z$ are a natural generalization of the qubit Pauli operators $\sigma_{x}$ and $\sigma_{z}$ as they act as $X|k\rangle=|(k+1)\,\mathrm{mod}\,d\,\rangle$ and $Z|k\rangle=\omega^{k}|k\rangle$ where $\omega=\exp(2\pi\mathrm{i}/d)$ is the first $d^{\text{th}}$ root of unity. They obey the relation $Z^{z}X^{x}=\omega^{xz}X^{x}Z^{z}$ for $x,z\in\mathbb{Z}_{d}$ and are unitary with $X^{-1}=X^{d-1}=X^{\dagger}$ and $Z^{-1}=Z^{d-1}=Z^{\dagger}$. For the general qudit Pauli operators, we use the Heisenberg-Weyl displacement operators
$T_{(x,z)}=\omega^{2^{-1}xz}X^{x}Z^{z}\,.$ (2)
The qudit Pauli group $\mathcal{P}$ can be decomposed into $d+1$ subgroups of commuting operators $\mathcal{G}_{r}=\{T_{(s,sr)}\,,s\in\mathbb{Z}_{d}\}\subset\mathcal{P}$ for $r=0,\dots,d-1$ and $\mathcal{G}_{d}=\{T_{(0,z)}\,,z\in\mathbb{Z}_{d}\}$. The eigenbases of two elements of distinct groups are mutually unbiased, i.e., their respective eigenvector’s inner product is of magnitude $1/\sqrt{d}$. Taking an element from each commutative group defines a complete set of mutually unbiased bases through their eigenvectors. Any qudit state has a unique representation $\rho=\frac{1}{d}\sum_{u\in\mathbb{Z}_{d}^{2}}\chi_{u}(\rho)T_{u}$ via its complex characteristic function
$\chi_{u}(\rho)=\mathrm{Tr}(\rho T_{-u})\,.$ (3)
The set of Hermitian operators obtained by applying the Fourier transform to Pauli operators forms an orthonormal basis in the vector space of operators equipped with the Hilbert-Schmidt inner product $(A,B):=d^{-1}\mathrm{Tr}(A^{\dagger}B)$. Writing $u=(u_{x},u_{z})\in\mathbb{Z}_{d}^{2}$ and using the symplectic form $[u,v]:=u_{z}v_{x}-u_{x}v_{z}$, the operators
$A_{u}=\frac{1}{d}\sum_{v\in\mathbb{Z}_{d}^{2}}\omega^{[u,v]}T_{v}\,,$ (4)
generate Gross’s Wigner function, $W_{u}(\rho):=d^{-1}\mathrm{Tr}\left(A_{u}\rho\right)$, a quasi-probability distribution that is real-valued and normalized, but can have negative values Gross2006 . Any qudit state can be written as $\rho=\sum_{u\in\mathbb{Z}_{d}^{2}}W_{u}(\rho)A_{v}$, and
$W_{u}(\rho)=\frac{1}{d}\sum_{v\in\mathbb{Z}_{d}^{2}}\omega^{-[u,v]}\chi_{v}(\rho)\,,$ (5)
For a system of $n$ qudits, we denote $\bm{u}=({\bm{u}_{x}},{\bm{u}_{z}})\in\mathbb{Z}_{d}^{2n}$, and the operator $T_{\bm{u}}=\bigotimes_{i=1}^{n}T_{((u_{i})_{x},(u_{i})_{z})}$, similarly for $A_{\bm{u}}$, which also define analogously the Wigner function $W_{\bm{u}}(\sigma)$ and characteristic function $\chi_{\bm{u}}(\sigma)$ for a state $\sigma$ on multiple qudits. In this context, we also use the inner product ${\bm{u}}\cdot{\bm{v}}=\sum_{i=1}^{n}u_{i}v_{i}$, which also extends to the symplectic form $[\bm{u},\bm{v}]=\bm{u}_{z}\bm{v}_{x}-\bm{u}_{x}\bm{v}_{z}$.
@@
Stabilizer states encompass a large class of states, including Bell states, GHZ states, and graph states. Each stabilizer state $|S\rangle$ is defined by an Abelian subgroup $\mathcal{S}$ of the Pauli group. We define the set ${\bm{u}}\in\Sigma\subset\mathbb{Z}_{d}^{2n}$ so that $\omega^{t({\bm{u}})}T_{\bm{u}}\in\mathcal{S}$ for a linear function $t$, which can be written as $t({\bm{u}})=[{\bm{a}},{\bm{u}}]$ for an ${\bm{a}}\in\mathbb{Z}_{d}^{2n}$. The stabilizer elements act trivially on the stabilizer state $\omega^{t({\bm{u}})}T_{{\bm{u}}}|S\rangle=|S\rangle$, while for any other $\omega^{t({\bm{v}})}T_{\bm{v}}\notin\mathcal{S}$, the commutation relation leads to $\langle S|T_{\bm{v}}|S\rangle=0$. The Wigner function of stabilizer states is non-negative and remains so under the action of Clifford operators, as their group is the normalizer of the Pauli group. Appleby appleby2005symmetric , Appleby2009 finds a unique representation for every Clifford operator with a decomposition into a Pauli operator and a unitary evolution under quadratic Hamiltonians.
Non-Clifford operations and Wigner negativity.
Howard and Vala Howard2012 provide a generalization of the qubit $T$-gate to qudits. They start from a diagonal unitary operator

### v4

# Qudit Clauser-Horne-Shimony-Holt Inequality and Nonlocality from Wigner Negativity

@@ Abstract
Nonlocality is an essential concept that distinguishes quantum from classical models and has been extensively studied in systems of qubits. For higher-dimensional systems, certain results for their two-level counterpart, like Bell violations with stabilizer states and Clifford operators, do not generalize. On the other hand, similar to continuous variable systems, Wigner negativity is necessary for nonlocality in qudit systems. We propose a new generalization of the CHSH inequality for qudits by inquiring correlations related to the Wigner negativity of stabilizer states under the adjoint action of a generalization of the qubit $\pi/8$-gate. A specified stabilizer state maximally violates the inequality among all qudit states based on its Wigner negativity. The Bell operator not only serves as a measure for the singlet fraction but also quantifies the volume of Wigner negativity. Additionally, we show how a bipartite entangled qudit state can serve as a witness for contextuality when it exhibits Wigner negativity. Furthermore, we give deterministic Bell violations, as well as violations with a constant number of measurements, for the Bell state relying on operators innate to higher-dimensional systems than the qudit at hand.
@@
Our main contribution is a novel generalization of the qubit Clauser-Horne-Shimony-Holt (CHSH) operator to qudits. This generalization is less restrictive than previous propositions allowing for extensions to the multipartite Bell scenario and can be tailored to a large class of entangled qudit states. There, we focus on qudit stabilizer states, which are uniquely defined by the Abelian subgroup of the qudit Pauli operators or Heisenberg-Weyl displacement operators, and encompass Bell states, GHZ states, and graph states. Contrary to qubits, Pauli operators on stabilizer states cannot invoke contextual violations —a prerequisite to nonlocality— due to their semi-classical nature exposed by a positive semi-definite Wigner function Kochen1975 , Howard2013 , Gross2006 . Similar to continuous variable systems PhysRevLett.129.230401 , the negativity of Gross’s Wigner function is equivalent to the contextuality of projective Pauli measurements in odd dimensions Howard2014 , Delfosse2017 . Therefore, Wigner-negativity, whether located in the state or the measurement, is a necessary requirement for nonlocality.
@@
For any odd prime dimension $d$, we first show in Eq. (9) how certain bipartite entangled states that exhibit Wigner negativity can be witnesses for nonlocality. Then, we construct a family of Bell operators on two qudits in Eq. (14) that relates to Gross’s Wigner function and provides a measure for the singlet fraction as well as for the volume of Wigner negativity. The Bell operators consist of qudit Pauli operators rotated by a generalization of the qubit $\pi/8$-gate introduced in Howard2012 , which we call unitary cube operators. Mapping the Pauli group to the qudit Clifford group appleby2005symmetric , Appleby2009 , unitary cube operators cause Wigner negativity in stabilizer states when these states are rotated by them. A corresponding Bell inequality is violated by the Bell state and is a natural extension of the CHSH inequality for qubits PhysRevLett.23.880 , which is similarly a linear combination of Pauli operators under the local rotation of the $\pi/8$-gate. We can then reduce the number of measurement settings constituting the Bell operator by only testing the stabilizer elements of the Bell state with the same implications as before. While the underlying mechanisms for Bell violations are equivalent to the ones Howard2015bell , our main contribution is a broader generalization of the qubit CHSH inequality for qudits adaptable to operators beyond the unitary cube operators, as well as a general insight into the origins of nonlocality in qudit systems. Lastly, this Bell operator can be adapted to any entangled qudit stabilizer state that maximally violates a corresponding Bell inequality, though it does not detect genuine multipartite entanglement beyond three parties. For $d>3$, unitary cube operators are inherently defined in the qudit system, that is, their eigenvalues are $d^{\text{th}}$ roots of unity, called characters of the finite field wiki:Dirichlet_character . Alternatively, we show strong nonlocal violations for Bell states in $d$ dimension with operators whose spectrum contains characters of yet higher-dimensional systems.
Settings.
Here, we restrict our study to odd prime dimensions, where all numbers are integers modulo $d$ forming a finite field, in particular, $2^{-1}$ is unambiguous. This also prevents studying properties inherent to subsystems, which is contiguous in systems of composite dimensions. Then, the qudit operators $X$ and $Z$ are a natural generalization of the qubit Pauli operators $\sigma_{x}$ and $\sigma_{z}$ as they act as $X|k\rangle=|(k+1)\,\mathrm{mod}\,d\,\rangle$ and $Z|k\rangle=\omega^{k}|k\rangle$ where $\omega=\exp(2\pi\mathrm{i}/d)$ is the first $d^{\text{th}}$ root of unity. They obey the relation $Z^{z}X^{x}=\omega^{xz}X^{x}Z^{z}$ for $x,z\in\mathbb{Z}_{d}$ and are unitary with $X^{-1}=X^{d-1}=X^{\dagger}$ and $Z^{-1}=Z^{d-1}=Z^{\dagger}$. For the general qudit Pauli operators, we use the Heisenberg-Weyl displacement operators
$T_{(x,z)}=\omega^{2^{-1}xz}X^{x}Z^{z}\,.$ (1)
The qudit Pauli group $\mathcal{P}$ can be decomposed into $d+1$ subgroups of commuting operators $\mathcal{G}_{r}=\{T_{(s,sr)}\,,s\in\mathbb{Z}_{d}\}\subset\mathcal{P}$ for $r=0,\dots,d-1$ and $\mathcal{G}_{d}=\{T_{(0,z)}\,,z\in\mathbb{Z}_{d}\}$. Any qudit state has a unique representation $\rho=\sum_{u\in\mathbb{Z}_{d}^{2}}\chi_{u}[\rho]T_{u}/d$ via
$\chi_{u}[\rho]=\mathrm{Tr}(\rho\,T_{-u})\,,$ (2)
which straightforwardly extends to any number of qudits, in particular $\chi_{u,v}[\rho]=\mathrm{Tr}(\rho\,T_{-u}\otimes T_{-v})$ if $\rho$ is a two-qudit state. The discrete Fourier transform of the characteristic function is Gross’s Wigner function, a quasi-probability distribution that is real-valued and normalized, but can have negative values Gross2006 . Writing $u=(u_{x},u_{z})\in\mathbb{Z}_{d}^{2}$ and using the symplectic form $[u,v]:=u_{z}v_{x}-u_{x}v_{z}$, it is
$W_{u}[\rho]=\frac{1}{d}\sum_{v\in\mathbb{Z}_{d}^{2}}\omega^{-[u,v]}\chi_{v}[\rho]\,.$ (3)
For a system of $n$ qudits, we denote $\bm{u}=({\bm{u}_{x}},{\bm{u}_{z}})\in\mathbb{Z}_{d}^{2n}$, and the operator $T_{\bm{u}}=\bigotimes_{i=1}^{n}T_{((u_{i})_{x},(u_{i})_{z})}$, similarly for $A_{\bm{u}}$, which also define analogously the Wigner function $W_{\bm{u}}[\rho]$ and characteristic function $\chi_{\bm{u}}[\rho]$ for a state $\rho$ on multiple qudits. In this context, we also use the inner product ${\bm{u}}\cdot{\bm{v}}=\sum_{i=1}^{n}u_{i}v_{i}$, which also extends to the symplectic form $[\bm{u},\bm{v}]=\bm{u}_{z}\bm{v}_{x}-\bm{u}_{x}\bm{v}_{z}$.
@@
Studying nonlocality in qudit systems, in a Bell scenario for $n$ parties, each party performs a local measurement $O_{m_{i}}$ depending on the local measurement setting $m_{i}$. Physical models can pose restrictions on the general probability $p({\bm{a}}|{\bm{m}})=p(\,(a_{i})_{i}\,|\,(m_{i})_{i}\,)$ of the outcome ${\bm{a}}$ for the measurement settings ${\bm{m}}$. In the quantum model, the probability distribution is derived from Born’s rule $p({\bm{a}}|{\bm{m}})=\mathrm{Tr}\left(\Pi^{{\bm{m}}}_{a^{\ast}}\rho\right)$ for a $n$-partite quantum state $\rho$ and projectors $\Pi^{{\bm{m}}}_{a^{\ast}}$ of the joint measurement $\bigotimes_{i=1}^{n}O_{m_{i}}$ with outcome $a^{\ast}=\prod_{i=1}^{n}a_{i}$. The (classical) local-hidden-variable (lhv) model is more restrictive with probability distributions of the form $p({\bm{a}}|{\bm{m}})=\sum_{\lambda}\mu(\lambda)\prod_{i=1}^{n}p_{i}(a_{i}|m_{i},\lambda)\,,$ for probability distributions $\mu$ and $p_{i}$. The separation can be proven with a Bell inequality for a Bell operator $\mathcal{B}=\sum_{{\bm{m}}}c_{{\bm{m}}}\bigotimes_{i=1}^{n}O_{m_{i}}\,,$ with coefficients $c_{{\bm{m}}}$. If an lhv model describes the probability distribution $p$, then
@@
and $\left\langle\mathcal{B}\right\rangle_{\mathrm{lhv}}\leq\max_{p\in\mathrm{lhv}}\left\langle\mathcal{B}\right\rangle_{\mathrm{lhv}}:=B_{\mathrm{lhv}}$ is a Bell inequality. Any distribution $p({\bm{a}}|{\bm{m}})$ such as $\sum_{{\bm{a}},{\bm{m}}}c_{{\bm{m}}}{\bm{a}}^{\ast}p({\bm{a}}|{\bm{m}})>B_{\mathrm{lhv}}$ defies classical local models. In this work, the measurements that define the Bell experiment are projective measurements with non-degenerate eigenvalues as outcomes.
A new qudit CHSH operator.
A notorious bipartite Bell inequality for experiments with two measurement outcomes is the CHSH inequality, which is violated by local Clifford measurement on two qubits in a Bell state. The corresponding Bell operator, written in terms of operators $B_{i}$ and $C_{i}$, is
$\mathcal{B}_{2}=\sum_{i,j=0,1}(-1)^{ij}B_{i}\otimes C_{j}\,.$ (5)
A natural generalization of this inequality to qudits considers the $d^{\text{th}}$ root of unity $\omega=\exp{(2\pi\mathrm{i}/d)}$, instead of the second root of unity $-1$, such that
$\mathcal{B}_{d}=\sum_{k,i,j=0}^{d-1}\omega^{nij}B_{i}^{k}\otimes C_{j}^{k}\,.$ (6)
The sum over $k$ ensures that the operator is Hermitian. This operator and related ones have been studied in several works Liang2009 , Ji2008 , salavrakos2017bell , Kaniewski2018 , Howard2015bell ,which could show nonlocality with local qudit Clifford measurement on two qudits in a qudit Bell state, or local Pauli operators and Bell states rotated by local non-Clifford operators.
Our main result stems from the simple observation that the qubit Bell operator (5) can also be written as
$\mathcal{B}_{2}=\sum_{u,v=(0,1),(1,0)}\chi_{u,v}[|\Phi_{\pi/8}\rangle\langle\Phi_{\pi/8}|]\,B_{u}\otimes C_{v}\,,$ (7)
with the characteristic function for bipartite qubit states
$\chi_{u,v}[\rho]=\mathrm{Tr}(\rho\,\sigma_{u}\otimes\sigma_{v})\,,$
the Pauli operator $\sigma_{(0,0)}=\mathds{1}$, $\sigma_{(0,1)}=\sigma_{z}$, $\sigma_{(1,0)}=\sigma_{x}$, $\sigma_{(1,1)}=\sigma_{y}$, and the rotated Bell state $|\Phi_{\pi/8}\rangle=\mathds{1}\otimes R_{\pi/8}\left(|00\rangle+|11\rangle\right)/\sqrt{2}\,,$ and the non-Clifford operator $R_{\pi/8}=\exp(\mathrm{i}\pi\sigma_{z}/8)$. Generalizing the CHSH operator (7) to $\mathcal{B}^{\prime}_{2}=\sum_{u,v=\mathbb{Z}_{2}^{2}}\chi_{u,v}[|\Phi_{\pi/8}\rangle\langle\Phi_{\pi/8}|]B_{u}\otimes C_{v}$ adds two non-zero terms that can be treated independently and do not change the nonlocal violation, see Appendix B.
Then, a natural generalization of this qubit Bell operator is the qudit Bell operator
$\mathcal{B}[\rho]=\sum_{u,v\in\mathbb{Z}_{d}^{2}}\chi_{u,v}[\rho]\,B_{u}\otimes C_{v}$ (8)
for some bipartite qudit state $\rho$, which might well be the rotated qudit Bell state $\rho=|\Phi_{U}\rangle\langle\Phi_{U}|$ with $|\Phi_{U}\rangle=\mathds{1}\otimes U\sum_{j\in Z_{d}}|jj\rangle/\sqrt{d}\,,$ for some unitary operator $U$.
Even though, they are both natural extensions of qubit Bell operator (5), the Bell operator (8) and the Bell operator in Eq. (6) are intrinsically different as Eq. (8) has $d^{2}$ inputs per side compared to $d$ in Eq. (6). With the right normalization, interpreting its coefficients $\omega^{nij}$ as entries of a characteristic function might be possible since it only defines $1+(d-1)d^{2}$ entries while the characteristic function of a bipartite state has $d^{4}$ entries. In contrast, Eq. (8) defines a large class of Bell operators for technically any bipartite qudit state. While most states do not allow the Bell operator (8) to witness nonlocality, we find that the qudit Bell state rotated by unitary non-Clifford operators defines a Bell inequality that the Bell state maximally violates among all qudit states.
In the following, we consider two equivalent quantum realizations for the Bell operator (8). One with a pure state defining the operator and Pauli measurements, and the other with the Bell state and Clifford measurements. For the later, we study extensions of the operator to one with fewer measurements and tailored to any qudit stabilizer state. In the first case, a specific state $\rho$ becomes, first, a witness for contextuality and, then, for nonlocality. The operator, corresponding to a entangled two-qudit state $\rho$ is
$\mathcal{B}[\rho]=\rho=\frac{1}{d^{2}}\sum_{u_{1},u_{2}\in\mathbb{Z}_{d}^{2}}\chi_{u_{1},u_{2}}[\rho]\,T_{u_{1}}\otimes T_{u_{2}}\,.$ (9)
The corresponding measurements are defined in terms of projectors $\Pi^{(v)}_{m}=d^{-1}\sum_{k\in\mathbb{Z}_{d}}\omega^{mk}T_{v}^{k}$ of outcome $m\in\mathbb{Z}_{d}$ and $v\in\mathbb{Z}_{d}^{2}$ on the eigenbasis of the Pauli operators. Although the operator comprises of $d^{4}$ local operators $T_{v}$, only $(d+1)^{2}$ measurements corresponding to mutually unbiased bases for each qudit are necessary. As such the Bell operator (9) might be seen as one leading to a $I(d+1)(d+1)dd$ inequality, a generalization of the $I3322$ inequality from Froissart1981 .
$\langle\mathcal{B}[\rho]\rangle_{\mathrm{nc}^{\ast}}$ $=\frac{1}{d^{2}}\sum_{u_{1},u_{2}\in\mathbb{Z}_{d}^{2}}\chi_{u_{1},u_{2}}[\rho]\,\omega^{[a_{1},{u_{1}}]}\omega^{[a_{2},{u_{2}}]}$
$=d^{2}\,W_{a_{1},a_{2}}[\rho]\,.$ (10)
$\left\langle\mathcal{B}[\rho]\right\rangle_{\mathrm{nc}}\leq d^{2}\max_{a_{1},a_{2}\in\mathbb{Z}_{d}^{2}}|W_{a_{1},a_{2}}[\rho]|\equiv B_{\mathrm{nc}}^{\max}\,.$ (11)
On the other hand, a pure state state $\rho$ maximally violates the non-contextual inequality, which it defines via its characteristic function, among all qudit states if $\mathrm{Tr}(\rho^{2})=1>B_{\mathrm{nc}}^{\max}$. We confirm with Eq. (36) that non-trivial negativity volume, $N[\Psi]>0$, is necessary for such a violation. As a result, Eq. (9) defines a family of inequalities based on Wigner negative states that are witnesses of contextuality if $1>d^{2}\max_{a_{1},a_{2}\in\mathbb{Z}_{d}^{2}}|W_{a_{1},a_{2}}[\rho]|$. This specifically exploit the equivalence of Wigner negativity and contextuality with Pauli measurements Delfosse2017 .
$\langle\mathcal{B}[\rho]\rangle_{\mathrm{lhv}}\leq\max_{\alpha,\beta\in\mathbb{Z}_{d}^{d+1}}\sum_{u_{1},u_{2}\in\mathbb{Z}_{d}^{2}}\chi_{u_{1},u_{2}}[\rho]\,\omega^{\alpha_{r(u_{1})}+\beta_{r(u_{2})}}\equiv B^{\max}_{\mathrm{lhv}}\,.$ (12)
In terms of the Wigner function, the Bell bound is
$B^{\max}_{\mathrm{lhv}}=\max_{\alpha,\beta\in\mathbb{Z}_{d}^{d+1}}\sum_{u_{1},u_{2}\in\mathbb{Z}_{d}^{2}}W_{u_{1},u_{2}}[\rho]\,F(\alpha)_{u_{1}}F(\beta)_{u_{1}}\,.$ (13)
with the discrete Fourier transform of $\omega^{a_{r(u)}}$, $F(a)_{u}=\sum_{v\in\mathbb{Z}_{d}^{2}}\omega^{[u,v]+\alpha_{r(v)}}/d$. The values of $F(\alpha)_{u}$ range between $-1,0,\dots,d$, the sum of entries is $\sum_{u\in\mathbb{Z}_{d}^{2}}F(\alpha)_{u}=d$, and the vector’s norm is equally $\|F(\alpha)\|_{2}=d$. For an $\alpha\in\mathbb{Z}_{d}^{d+1}$ such that $F(\alpha)_{v}=d$ for some $v\in\mathbb{Z}_{d}^{2}$, it is $\alpha_{r(u)}=[a,u]$ for some $a\in\mathbb{Z}_{d}^{2}$, and $F(\alpha)_{w}=0$ for all $w\in\mathbb{Z}_{d}^{2}\setminus\{v\}$, which is precisely, the non-contextual value assignment used for the Wigner negativity witness or non-contextual inequality (11). To evaluate the Bell bound (13) further, the Cauchy-Schwarz inequality gives $B^{\max}_{\mathrm{lhv}}[\rho]\leq\|W[\rho]\|_{2}\|F(\alpha)\|_{2}\|F(\beta)\|_{2}\leq d^{2}\|W[\rho]\|_{2}$. To observe a Bell violation, it is necessary that $B_{\mathrm{lhv}}[\rho]<1$, which requires at least $\|W[\rho]\|_{2}\leq 1$. Since the non-contextual value assignments are a special case of the deterministic lhv models’, it is $B^{\max}_{\mathrm{nc}}\leq B^{\max}_{\mathrm{lhv}}$ such that Wigner negativity of the pure resource state $\rho$ as well as the state used for the quantum realization is again essential for nonlocality. Whether the state defines a notable Bell inequality beyond these criteria requires an analysis of the state’s characteristic or Wigner function. Bell states rotated by local unitary non-Clifford operators are prime candidates to define a Bell inequality in terms of their characteristic function.
@@
Qudit stabilizer states, however, do not exhibit Wigner negativity. Stabilizer states encompass a large class of states, including Bell states, GHZ states, and graph states. Each stabilizer state $|S\rangle$ is defined by an Abelian subgroup $\mathcal{S}$ of the Pauli group. The Wigner function of stabilizer states is non-negative and remains so under the action of Clifford operators, as their group is the normalizer of the Pauli group. This motivates the quantum realization of the Bell operator (8) that is similar to the one for the qubit Bell operator (5) leading to a violation of the CHSH inequality with the Bell state. For the qudit Bell state the resulting operator is
$\mathcal{B}_{U}=\frac{1}{d^{2}}\sum_{u_{1},u_{2}\in\mathbb{Z}_{d}^{2}}\chi^{U}_{u_{1},u_{2}}~{}U\,T_{u_{1}}U^{\dagger}\otimes T_{u_{2}}\,,$ (14)
with the characteristic function $\chi^{U}_{u_{1},u_{2}}:=\langle\Phi_{U}|T_{-u_{1}}\otimes T_{-u_{2}})|\Phi_{U}\rangle$ of the rotated Bell state. The non-contextual bounds arise for the same reasons as those underlying the inequality (11) such that $\left\langle\mathcal{B}_{U}\right\rangle_{\mathrm{nc}}\leq d^{2}\max_{v_{1},v_{2}\in\mathbb{Z}_{d}^{2}}|W_{v_{1},v_{2}}^{U}|\equiv B_{\mathrm{nc}}^{\max}\,.$ The Bell operator’s expectation value for the Bell state is $\mathrm{Tr}\left(\mathcal{B}_{U}|\Phi\rangle\langle\Phi|\right)=\sum_{v_{1},v_{2}\in\mathbb{Z}_{d}^{2}}d^{2}(W_{v_{1},v_{2}}^{U})^{2}=1\,,$ since the rotated Bell state is a pure state. Moreover, we show in Eq. (31) that the expectation value of the Bell operator for any bipartite state $\rho$ is a measure for the singlet fraction
$\langle\mathcal{B}_{U}\rangle_{\rho}=\mathrm{Tr}\left(\mathcal{B}_{U}\rho\right)=\langle\Phi|\rho|\Phi\rangle\,.$ (15)
From Eq. (37), the expectation value with a quantum state $\rho$ can be bound by $B_{\mathrm{nc}}^{\max}$ and the volume of Wigner negativity using Hölder’s inequality with the 1-norm and the maximum norm,
$\mathrm{Tr}\left(\mathcal{B}_{U}\rho\right)\leq B_{\mathrm{nc}}^{\max}(1+2N[\rho^{\prime}])\,,$ (16)
with $\rho^{\prime}:=(U\otimes\mathds{1})^{\dagger}\rho(U\otimes\mathds{1})$. In particular, it is $1\leq B_{\mathrm{nc}}^{\max}(1+2N[|\Phi_{U}\rangle\langle\Phi_{U}|])\,.$ Thus, a non-contextual violation $\mathrm{Tr}\left(\mathcal{B}_{U}|\Phi\rangle\langle\Phi|\right)=1>B_{\mathrm{lhv}}^{\max}$ is possible for a nontrivial negativity volume, $N[|\Phi_{U}\rangle\langle\Phi_{U}|]>0$.
As an example, we consider the generalization of the qubit $\pi/8$-gate to a qudit one that was introduced by Howard and Vala Howard2012 and also used by Howard in Howard2015bell to show nonlocality with a Bell operator similar to the one in Eq. (6). The authors start from a diagonal unitary operator